OpenStudy (anonymous):

Solve. 1/3x+9-2/x+3=2

OpenStudy (anonymous):

Where is the problem?

OpenStudy (ledah):

I believe that is the problem

OpenStudy (anonymous):

Yes, but we need to know why she has doubts to solving it.

OpenStudy (anonymous):

\[1\div3x+9-2\div x+3=2\] (the divison signs are supposed to be long like : ----

OpenStudy (anonymous):

At what part you get stucked?

OpenStudy (anonymous):

oh I don't understand how to solve it

OpenStudy (anonymous):

\[1/(3x+9)-2/(x+3)=2\] Is it correct like this?

OpenStudy (ledah):

If the problem should look like marrioncito's then you get 1/2 as the answer.

OpenStudy (anonymous):

Yes Marrioncito and thank you Ledah :)

OpenStudy (ledah):

Your welcome :)

OpenStudy (anonymous):

\[1/(3x+9) =2 + 2/(x+3)\]

OpenStudy (anonymous):

wait a second...... would the answer change if there were no parenthses, cause there are none in the problem

OpenStudy (anonymous):

You got a different problem.

OpenStudy (anonymous):

\[1/3x+9−2/x+3=2\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[1/3x - 2/x = 2 -3 - 9 = -10\]

OpenStudy (anonymous):

\[1/3x - 2/x = -5/3x = -10\] \[5/3x=10\] \[5=10.3x = 30x\] \[5 = 30x\] \[5/30 = x\] \[1/6 = x\]

OpenStudy (anonymous):

Leda has not read that the problem was another one. Follow my steps, if you have questions just ask!

OpenStudy (anonymous):

here I'm not sure you guys are getting the problem....let me draw it for you

OpenStudy (anonymous):

I rewritted it like this: to Medals 0 1/3x+9−2/x+3=2

OpenStudy (anonymous):